# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

### v1

@@ 4.2. Quantization
Overall, quantization is one of the easiest GNN acceleration methods to apply but shows significant speed improvement. It can also be combined with other acceleration methods to further boost speed gain. For example, (Zhao et al., 2020) combines quantization with network architecture search (NAS). In (Zhao et al., 2020), a quantization search space are specified and the best quantized GNN is selected via NAS.
4.3. Distillation
Knowledge distillation (KD) or simply distillation is a technique to compress classification models to smaller and thus faster models (Buciluǎ et al., 2006). It was shown to be able to produce much faster models with very little, if not none, accuracy loss for image classification and speech recognition (Hinton et al., 2015). The basic idea of KD is to train two models in a teacher-student fashion. The teacher outputs a probability vector for each input data containing probabilities for the data to be classified into $C$ classes. The probability vectors (or a re-normalized version of them) are in $[0,1]^{C}$ and are called soft labels, in contrast to the ground-truth hard labels in $\{0,1\}^{C}$. Such soft labels will be used to train a simpler student model. Specifically, the standard KD loss on a data instance $i$ is computed as a distance between the teacher model output (soft label) of data $i$, i.e., $\hat{\bm{y}}_{i}^{t}$, and the student model output of data $i$, i.e., $\hat{\bm{y}}_{i}^{s}$. Two outputs are matched by minimizing this loss, and thus knowledge is transferred from the teacher to the student. Since both outputs are probability vectors, a common choice of the distance metric is the Kullback-Leibler (KL) divergence denoted as $D_{KL}$,
(14) $D_{KL}(\hat{\bm{y}}_{i}^{t}||\hat{\bm{y}}_{i}^{s})=\sum_{j}\hat{\bm{y}}_{ij}^{t}\log\frac{\hat{\bm{y}}_{ij}^{t}}{\hat{\bm{y}}_{ij}^{s}}.$
In most cases, the teacher model can be trained alone first, and the student is trained next with the trained teacher fixed. Thus, we can drop the constant $\hat{\bm{y}}_{ij}^{t}\log\hat{\bm{y}}_{ij}^{t}$ term to get the form after $\propto$ for student training. The knowledge distillation loss $\mathcal{L}_{i}^{KD}$ will only be used to update the parameters of the student but not the teacher,
(15) $\mathcal{L}_{i}^{KD}\propto D_{KL}(\hat{\bm{y}}_{i}^{t}||\hat{\bm{y}}_{i}^{s})=-\sum_{j}\hat{\bm{y}}_{ij}^{t}\log\hat{\bm{y}}_{ij}^{s}.$
KD can be naturally applied in a semi-supervised setting, a common setting for graph node classification, by training the teacher with labeled data and generating soft labels for the unlabeled data to train the student (Xie et al., 2020). KD has been wildly used for compressing and accelerating DNNs like CNNs (Hinton et al., 2015), and recently it has been extended to GNNs as well (Yang et al., 2020; Yan et al., 2020b; Zhang et al., 2022a). We discuss these GNN KD methods in detail below. They mostly follow the KD framework discussed above and mainly differ in two perspectives: 1) What are the teacher and the student models? 2) What is the KD objective?
LSP (Yang et al., 2020) is the pioneer work exploring KD for accelerating GNNs. Targeting node classification, the teacher model LSP considers is a GAT and the student model is a GAT with fewer parameters. LSP proposes a new KD objective termed local structure preserving loss. The intuition is that node representations are generated by aggregating messages within the local structure, so the desired distillation function should transfer not only the final aggregation result but also the local structure as part of the knowledge. LSP defines the local structure as the similarity between each node to its neighbors and encourages the local structure of the student to match the teacher. Specifically, the local structure of node $i$ is represented as a vector ${\bm{s}}^{(i)}\in{\mathcal{R}}^{d(v)}$, with ${\bm{s}}^{(i)}_{j}$ representing the local influence of node $j$ to node $i$ defined as
(16) ${\bm{s}}^{(i)}_{j}=\frac{\exp(\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{j}))}{\sum_{k:(k,i)\in{\mathcal{E}}}\exp(\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{k}))},\text{ where }\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{k})=||{\bm{h}}_{i}-{\bm{h}}_{k}||_{2}^{2}.$
Plug-in node representations ${\bm{h}}$ of the teacher or the student gives the teacher local structure ${\bm{s}}^{(i),t}$ and the student local structure ${\bm{s}}^{(i),s}$ respectively. Then, the local structure preserving loss $\mathcal{L}_{i}^{LSP}$ is defined using the KL divergence, similarly as the $\mathcal{L}_{i}^{KD}$ in Equation (17). The difference being $\mathcal{L}_{i}^{LSP}$ computes the KL-divergence of ${\bm{s}}^{(i),s}$ against ${\bm{s}}^{(i),t}$, instead of the teacher probability output $\hat{\bm{y}}_{i}^{t}$ against student output $\hat{\bm{y}}_{i}^{s}$. That is,
@@ 4.3. Distillation
TinyGNN (Yan et al., 2020b) shares a similar intuition as LSP to preserve local structures, but it chooses to add a special layer termed a Peer-Aware Module (PAM) to the student model instead of defining a new KD objective as in LSP. TinyGNN adopts the standard KL divergence loss $\mathcal{L}_{i}^{KD}$. The teacher model TinyGNN considers is GAT. The student model, however, is a GAT with fewer message-passing layers than the teacher but a special PAM layer inserted before each message-passing layer. The idea of PAM is to encourage nodes in the same message-passing hierarchy to directly interact. For example, when the GNN is only one layer, a message passing step centered at node $v$ will aggregate messages from its neighbors ${\mathcal{N}}(v)$, but nodes in ${\mathcal{N}}(v)$ will not directly interact with each other during this message passing step. In other words, the computation graph is a one-layer tree and has no edges between its leaves. TinyGNN argues that sharing messages between nodes in ${\mathcal{N}}(v)$ is important for capturing the local structure around $v$. This message sharing happens naturally when the GNN is deep like the teacher, but it is lacking if the GNN is only one or two layers like the student in TinyGNN. Thus, the PAM is proposed to encourage this message sharing by adding a self-attention layer between all nodes in ${\mathcal{N}}(v)$. It was shown that a GAT student with PAM has a higher distilled model accuracy than a regular GAT student, though the speed gain is not as much given the extra self-attention computation.
GLNN (Zhang et al., 2022a) explores KD for accelerating GNNs by making the student model as simple as a “graph-less neural network”, i.e., an MLP. GLNN follows the standard KL divergence loss $\mathcal{L}_{i}^{KD}$, but the student model in GLNN is a pure MLP (a zero-layer GNN) with no node dependency. GLNN shows that even with a student as simple as an MLP, KD can help the student MLP to achieve competitive performance to the teacher GNN on many benchmark datasets, even when used for the inductive prediction on unseen new nodes. GLNN can accelerate GNNs to orders of magnitude faster because the node dependency is dropped for the student. It is also shown to work with different teacher GNNs including GCN, GAT, GraphSAGE, etc. However, GLNN is still not a universal solution for all cases. The analysis in (Zhang et al., 2022a) shows that GLNN fails to work when the conditional mutual information between the node features ${\bm{X}}$ and the node labels ${\bm{Y}}$ given the graph structure ${\bm{A}}$ is low. Intuitively, if node labels depend mainly on the graph structure but not the node feature, then MLPs won’t be able to classify the nodes even with KD.
Overall, the methods above apply the general KD idea to GNNs by adapting graph properties different perspectives, including both new objects and new student designs. They all show exciting improvements over the original teacher GNN, but this line of research is still at an early stage. These KD methods were developed under different settings. Thus, we are unaware of the applicability of these distillation techniques to each others setting and whether we can combine them. Since KD in general is always a trade-off between model accuracy and model speed, we don’t know which method or which combination will give the best result balancing these two perspectives. Also, we don’t know under what exact conditions, simple methods like GLNN (Zhang et al., 2022a) will be applicable. These many interesting open questions make KD a promising direction to explore.
@@
Finally, there are some other related works applying KD techniques on GNNs, for which the goal is less accuracy loss but not necessarily inference acceleration (Yang et al., 2021b; Chen et al., 2021a; Zheng et al., 2021a). Different from the methods we introduced above, these works often choose students that are not necessarily simpler than the teacher, which can sometimes outperform the teacher in terms of accuracy, but they could be even slower than the teacher. We only briefly mention them here given they are less related to acceleration. GNN Self-Distillation (GNN-SD) (Chen et al., 2021a) considers teacher-free self-distillation where the knowledge is extracted and transferred between layers of the same GNN. GNN-SD is shown to consistently achieve better accuracy than the standard KD objective $\mathcal{L}_{i}^{KD}$. Combination of Parameterized label propagation and Feature transformation (CPF) (Yang et al., 2021b) combines KD and label propagation (LP) together. The student in CPF includes an MLP module for learning feature information and an LP module for learning structure information. The classification accuracy of a CPF student can generally outperform the accuracy of teacher GNNs after KD. However, since the LP module still has node dependency like the message passing (labels instead of the hidden features are passed as messages), the speed gain is likely limited and not discussed in (Yang et al., 2021b). Cold Brew (Zheng et al., 2021a) utilizes KD to improve GNN performance on nodes with incomplete or missing neighbors, i.e., the cold-start nodes. Cold Brew greatly improves the accuracy on these cold-start nodes, but its inference stage requires a scan through the whole graph to select existing nodes as neighbor candidates for the cold-start nodes. Cold Brew thus has even more node dependency and runs slower than a standard message-passing GNN. More related works include Graph-Free Knowledge Distillation (GFKD) (Deng and Zhang, 2021) transfers knowledge from a teacher GNN to a student GNN via generating fake graphs. GFKD focused on graph-level prediction tasks, for which data instances are independent graphs and the scalability problem caused by node dependence is less urgent.
4.4. Combine Together
Inference acceleration techniques in this section have been combined together to achieve better results. For example, quantization is combined with distillation in (Bahri et al., 2021). In (Bahri et al., 2021), GNNs weights and activations are excessively quantized to binary. This quantization is very ambitious and hard to achieve in one-step. (Bahri et al., 2021) uses a cascaded distillation scheme to gradually quantize more and more parts of the GNN and distill knowledge step-by-step to the next more quantized GNN down the sequence. The final quantized-and-distilled binary GNN can achieve 2X speedup on Raspberry Pi 4B with only moderate accuracy loss.
5. GNN Acceleration: COTS Systems
In addition to the efficient algorithms for GNN training/inference, optimizing the underlying system is essential for improving end-to-end throughputs of GNNs. In particular, existing works accelerate GNN systems from three perspectives: GPU kernel acceleration, user-defined function (UDF) optimization, and scalable training systems. Different from most deep learning (DL) tasks, GNN computation follows a message-passing paradigm where it is highly desirable to have efficient sparse operations; however, this cannot be sufficiently handled by existing DL platforms and infrastructure. Various efficient GNN kernels and UDF optimization techniques have been proposed for bridging the gap between the lack of fundamental support and high demand. Furthermore, due to the irregular dependency among nodes, training a graph with a huge number of nodes and a giant feature matrix is particularly challenging. How to design an efficient scalable GNN training system is still an open research question for the ML system community. Currently, two widely used frameworks for GNNs are PyG (Fey and Lenssen, 2019) and DGL (Wang et al., 2019b).
5.1. GPU Kernel Acceleration
GPUs are powerful hardware accelerators that can greatly speed up computations required in DNN training and inference. Under this context, GPU kernels refer to specialized programs optimized to be executed in parallel on the many processing units of a GPU, enabling DNN computations to be much faster and more efficient. Although GPUs are popular in being used for DL acceleration, accelerating GNNs with GPUs is still challenging due to the unique sparsity and irregularity in graphs. In recent years, many works propose novel GPU kernels for accelerating GNN workloads. Table 3 summarizes recent works targeting different GNN kernels as detailed below.
Table 3. Summary of GPU Kernel Acceleration Methods.
Work Methods Benefits
PCGCN (Tian et al., 2020) Workload Reorder, Dual-Mode Computation Better Data Locality, Coalesced Memory Access
fuseGNN (Chen et al., 2020b) Kernel Fusion, Workload Reorder, Dual-Mode Computation Decreased Memory Traffic, Saved Memory
FusedMM (Rahman et al., 2021) Kernel Fusion Balanced Workload, Saved Memory
TLPGNN (Fu et al., 2022) Workload Reorder, Kernel Fusion, Dynamic Workload Assignment Decreased Memory Traffic, Coalesced Memory Access, Balanced Workload, No Branch Divergence
Zhang et al. (Zhang et al., 2022b) Workload Reorder, Kernel Fusion, Recomputing Embedding Reduced Computation, Reduced I/O, Saved Memory
@@ 5.1. GPU Kernel Acceleration
fuseGNN (Chen et al., 2020b) develops two workload abstractions of GNN aggregation, assuming that the input graph adopts the COO format: Gather-ApplyEdge-Scatter (GAS) and Gather-ApplyEdge-Reduce (GAR). Specifically, GAS works with the COO format graph, and processes edge-wise aggregation; GAR handles node-wise aggregation, which is particularly useful for the CSR format as we defined in Section 2.1. Although GAR is usually efficient because it relies on on-chip reduction with a shared output memory, creating CSR from COO is usually time costly, thus degrading the achievable efficiency of GNN computation. As a result, fuseGNN uses GAS for graphs with a low average degree (e.g., Cora(Sen et al., 2008a)) and GAR for graphs with a high average degree (e.g., Reddit (Hamilton et al., 2017a)). All operations in the three kernels, including graph processing for graph format conversion, GAS, and GAR, are fused for alleviating the memory traffic. fuseGNN is developed for PyTorch and can be extended to multi-GPU scenarios.
FusedMM (Rahman et al., 2021) is a CPU kernel but the authors claim that it can be extended to GPU acceleration. The work divides popular graph operators into five types of basic operations that can be launched sequentially (e.g., element-wise additive, reduction operation, element-wise self-operation) and defined by users for designing customized GNN layers. During GNN computations, the operations associated with each target node are fused together for achieving memory reduction by eliminating the necessity of saving intermediate results. Since each CPU thread is assigned multiple target nodes, FusedMM further adopts an edge-balanced partition to ensure a balanced workload among threads. The proposed kernel of FusedMM can be integrated into DGL (Wang et al., 2019b).
@@
Zhang et al. (Zhang et al., 2022b) first decompose a GNN operation into four parts (i.e., scatter across edges, apply edge, gather from edges, and apply node), and then identify three challenges in GNN computations: redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. Specifically, redundant neural operators can be caused when the same feature transformation is employed to the same node features after they are scattered across edges. To address this issue, the scatter operation can be postponed after the feature transformation is finished. For the second issue, if the thread mappings are inconsistent (e.g., scatter threads are assigned with edge-centric tasks, while gather threads are assigned with node-centric tasks), these threads have to be processed sequentially and cannot be merged. This work leverages unified thread mapping for fusion. The third issue identifies that intermediate results may occupy up to 92% of the total memory requirement, which can be resolved by recomputation during backward propagation.
Overall, workload reordering is necessary for identifying GPU-friendly processing patterns. In particular, most works sequentially compute output nodes and process features in parallel. The former allows kernel fusion and avoids branch divergence while the latter guarantees coalesced memory accesses. Because the graph structure can be highly irregular and can be presented in different formats, some work leverage dual-mode computation and dynamic workload assignment for handling graphs of different properties.
5.2. User-defined Function Optimization
The flexibility in selecting the message, aggregation, and update functions has been a critical component in the success of GNNs under the message-passing paradigm. GNN systems possess the ability for users to natively express these functions using tensor operators in Python and execute them through APIs provided by the system. This programming paradigm is termed user-defined functions (UDFs). Once defined, these functions can be reused throughout the program, increasing code modularity and reducing code duplication. Overall, UDFs provide a powerful tool for creating more efficient and maintainable codes. A few recent works have been developed to ease the deployment of user-defined GNNs.
FeatGraph (Hu et al., 2020c) provides a flexible programming interface on top of the compiler stack for DL systems: TVM (Chen et al., 2018b) for supporting friendly user-defined functions. The users can manually select the tiling factor (the number of features being processed in each thread) for maximizing the cache utilization. However, because TVM lacks the support for sparse operations, FeatGraph only supports SpMM/SDDMM-based operations.
@@ 5.2. User-defined Function Optimization
Graphiler (Xie et al., 2022) develops an automated tool for scatter/gather-related tasks. Specifically, compared with (Zhang et al., 2022b) mentioned in the previous section that manually implements operation reordering for some specific GNN operations, Graphiler enjoys better flexibility by automatically detecting redundant communication and then adjusting their order accordingly, and further fuses all operations after gathering from edges to reduce intermediate memory accesses.
Existing automation tools for GNN computation are not prosperous. This is because concurrent DL backends do not well support sparse operations. We believe this direction is getting mature as several recent backends for sparse DL operations have been proposed. For example, SparseTIR (Ye et al., 2022) is a general compiler infrastructure for sparse DL operations. It develops a set of intermediate representations that can represent operators supported by the aforementioned Seastar, FeatGraph, and Graphiler work.
5.3. Scalable Training Systems
To support the training of GNNs with large-scale graphs, many recent works have developed GNN systems targeting scalable training. From the system perspective, there are two types of scalability:
  • Vertical scalability handles scaling data with restricted computation resources. As the training graph of a GNN can be giant, this direction is highly practical. For the GNN systems pursuing vertical scalability, they usually assume the use of low-cost or unlimited host memory.
  • Horizontal scalability expects better efficiency with scaling resources. However, horizontal scalability is particularly challenging in GNN training because more workers incur more dependency across devices, leading to more significant communication and synchronization overhead and thus degrading the achievable training throughput.
Some concurrent works summarize recent distributed GNN training systems (Shao et al., 2022; Lin et al., 2022). In particular, Shao et al. (2022) classifies recent works in terms of hardware facilities (e.g., multi-GPU, GPU cluster, and CPU cluster). On the other hand, Lin et al. (2022) categorizes all works as full-batch training or mini-batch training. In this work, we focus on scalable GNN training, which is a wider topic than distributed GNN training, as it not only includes efficient distributed training (i.e., horizontal scalability) but also considers large-scale training with limited computing resources (i.e., vertical scalability). Furthermore, we observe that the key contribution of each system overlapped among (between) the categories adopted in existing survey papers, hindering the understanding of their technical differences. To facilitate a clearer comparison of the technical contributions made by different scalable GNN training systems, we categorize them based on the method by which data is loaded onto the computation devices: on-device systems, swap-based systems, and sampling-based systems. An overview of these categories is demonstrated in Figure 4, and we compare their scope in Table 4.
System Type Main Focus Challenges and Contribution
Swap-based system Full-graph aggregation Workload scheduler
Here we briefly summarize these three types of systems:
  • On-device Systems (Figure 4(a)) distributedly store the entire graph and associated features on the computation devices (e.g., GPUs or CPUs). Since each worker only maintains a portion of data, systems based on this storage architecture mostly focus on full-graph aggregation for better leveraging the computation resources. During training, each worker needs to transfer the intermediate embeddings for supporting the full-graph aggregation in the next layer. A naive system that follows this direction suffers from frequent and overwhelming communications as well as exploded memory requirements for caching the embeddings from other workers. Therefore, existing systems of this category mainly focus on maximizing communication and memory savings to enhance horizontal scalability.
  • Swap-based Systems (Figure 4(b)) store all data in (distributed) shared memory and swap it and the intermediate results to workers for computations. This approach enables each node to fully utilize its neighbors’ embeddings. However, developing a workload scheduler that ensures vertical scalability poses two challenges for swap-based systems: 1) balancing workload among different GPUs is non-trivial, and 2) naive full-graph aggregation with limited devices incurs long iteration latency because it only updates the model once per training epoch. Despite these challenges, effective solutions to this problem are crucial for improving the achievable efficiency of these systems and enhancing their overall performance.
  • Sampling-based Systems (Figure 4(c)) also save the whole graph and features in (distributed) shared memory; however, instead of performing full-neighbor aggregation, they adopt neighbor sampling for neighbor aggregation. To implement this approach, each sampler first determines sampled input nodes based on the output nodes for mini-batch training. The features of sampled input nodes are then fetched from the (distributed) shared memory. The primary challenge for sampling-based systems is data movement for mini-batch preparation, which can significantly limit performance. To overcome this bottleneck, these systems focus on developing better data movement and caching strategies to optimize communication. Overall, sampling-based systems enhance both horizontal and vertical scalability, making them a promising solution for GNN training on large-scale graphs.
(a) On-device Systems
(b) Swap-based Systems
(c) Sampling-based Systems
Figure 4. Overview of three types of scalable GNN training systems. The dashed lines in (a) indicate the dependency between nodes from different workers, while the grey nodes in (b) serve only as input nodes for subgraph computations. Model synchronization among workers is not shown for better visual clarity.
Table 5. Summary of Scalable Training Systems.
On-device System CAGNET (Tripathy et al., 2020) GPU Full-graph
Dorylus (Thorpe et al., 2021) Hybrid Full-graph
$P^{3}$ (Gandhi and Iyer, 2021) GPU Sampling
PipeGCN (Wan et al., 2022b) GPU Full-graph
LLCG (Ramezani et al., 2022) GPU Full-graph
BNS-GCN (Wan et al., 2022a) GPU Hybrid
SAR (Mostafa, 2022) CPU Full-graph
Sancus (Peng et al., 2022) GPU Full-graph
Swap-based System NeuGraph (Ma et al., 2019b) GPU Full-graph
Roc (Jia et al., 2020a) GPU Full-graph
GNNAdvisor (Wang et al., 2021a) GPU Full-graph
GNNAutoScale (Fey et al., 2021) GPU Full-graph
GraphFM (Yu et al., 2022) GPU Full-graph
Sampling-based System AliGraph (Zhu et al., 2019) CPU Sampling
PaGraph (Lin et al., 2020) GPU Sampling
DistDGL (Zheng et al., 2020a) CPU Sampling
DistDGLv2 (Zheng et al., 2021b) GPU Sampling
SALIENT (Kaler et al., 2022) GPU Sampling
GNNLab (Yang et al., 2022) GPU Sampling
BGL (Liu et al., 2021a) GPU Sampling
GNS (Dong et al., 2021) GPU Sampling
MariusGNN (Waleffe et al., 2022) GPU Sampling
5.3.1. On-device Systems
One convincing approach for GNN training with large-scale graphs is to partition the data into multiple workers. All workers update the local model using the locally labeled nodes and then synchronize the local models. This design is very similar to distributed data parallelism in DNN training. However, there are two fundamental differences: 1) GNN training requires feature transfer for each layer, which can explode the required communication volume; and 2) the node feature/embedding tensors are substantially larger than the model. These unique communication and memory challenges motivate the development of various GNN training systems. Since systems under this paradigm update models with dependent partitions instead of independent data, this approach is also named as partition parallelism.

### v2

@@ C.2. Quantization
C.3. Distillation
@@
TinyGNN (Yan et al., 2020a) shares a similar intuition as LSP to preserve local structures, but it chooses to add a special layer termed a Peer-Aware Module (PAM) to the student model instead of defining a new KD objective as in LSP. TinyGNN adopts the standard KL divergence loss $\mathcal{L}_{i}^{KD}$. The teacher model TinyGNN considers is GAT. The student model, however, is a GAT with fewer message-passing layers than the teacher but a special PAM layer inserted before each message-passing layer. The idea of PAM is to encourage nodes in the same message-passing hierarchy to directly interact. For example, when the GNN is only one layer, a message passing step centered at node $v$ will aggregate messages from its neighbors ${\mathcal{N}}(v)$, but nodes in ${\mathcal{N}}(v)$ will not directly interact with each other during this message passing step. In other words, the computation graph is a one-layer tree and has no edges between its leaves. TinyGNN argues that sharing messages between nodes in ${\mathcal{N}}(v)$ is important for capturing the local structure around $v$. This message sharing happens naturally when the GNN is deep like the teacher, but it is lacking if the GNN is only one or two layers like the student in TinyGNN. Thus, the PAM is proposed to encourage this message sharing by adding a self-attention layer between all nodes in ${\mathcal{N}}(v)$. It was shown that a GAT student with PAM has a higher distilled model accuracy than a regular GAT student, though the speed gain is not as much given the extra self-attention computation.
@@
We briefly mention more KD on graphs works and note that they are less related to acceleration. GNN Self-Distillation (GNN-SD) (Chen et al., 2021c) considers teacher-free self-distillation where the knowledge is extracted and transferred between layers of the same GNN. GNN-SD is shown to consistently achieve better accuracy than the standard KD objective $\mathcal{L}_{i}^{KD}$. Combination of Parameterized label propagation and Feature transformation (CPF) (Yang et al., 2021a) combines KD and label propagation (LP) together. The student in CPF includes an MLP module for learning feature information and an LP module for learning structure information. The classification accuracy of a CPF student can generally outperform the accuracy of teacher GNNs after KD. However, since the LP module still has node dependency like the message passing (labels instead of the hidden features are passed as messages), the speed gain is likely limited and not discussed in (Yang et al., 2021a). Cold Brew (Zheng et al., 2021b) utilizes KD to improve GNN performance on nodes with incomplete or missing neighbors, i.e., the cold-start nodes. Cold Brew greatly improves the accuracy on these cold-start nodes, but its inference stage requires a scan through the whole graph to select existing nodes as neighbor candidates for the cold-start nodes. Cold Brew thus has even more node dependency and runs slower than a standard message-passing GNN. More related works include Graph-Free Knowledge Distillation (GFKD) (Deng and Zhang, 2021) transfers knowledge from a teacher GNN to a student GNN via generating fake graphs. GFKD focused on graph-level prediction tasks, for which data instances are independent graphs and the scalability problem caused by node dependence is less urgent.
Appendix D Detailed Discussion of COTS Systems
D.1. GPU Kernel Acceleration
@@
fuseGNN (Chen et al., 2020b) develops two workload abstractions of GNN aggregation, assuming that the input graph adopts the COO format: Gather-ApplyEdge-Scatter (GAS) and Gather-ApplyEdge-Reduce (GAR). Specifically, GAS works with the COO format graph, and processes edge-wise aggregation; GAR handles node-wise aggregation, which is particularly useful for the CSR format as we defined in Section A.1. Although GAR is usually efficient because it relies on on-chip reduction with a shared output memory, creating CSR from COO is usually time costly, thus degrading the achievable efficiency of GNN computation. As a result, fuseGNN uses GAS for graphs with a low average degree (e.g., Cora(Sen et al., 2008)) and GAR for graphs with a high average degree (e.g., Reddit (Hamilton et al., 2017b)). All operations in the three kernels, including graph processing for graph format conversion, GAS, and GAR, are fused for alleviating the memory traffic. fuseGNN is developed for PyTorch and can be extended to multi-GPU scenarios.
FusedMM (Rahman et al., 2021) is a CPU kernel but the authors claim that it can be extended to GPU acceleration. The work divides popular graph operators into five types of basic operations that can be launched sequentially (e.g., element-wise additive, reduction operation, element-wise self-operation) and defined by users for designing customized GNN layers. During GNN computations, the operations associated with each target node are fused together for achieving memory reduction by eliminating the necessity of saving intermediate results. Since each CPU thread is assigned multiple target nodes, FusedMM further adopts an edge-balanced partition to ensure a balanced workload among threads. The proposed kernel of FusedMM can be integrated into DGL (Wang et al., 2019a).
@@
Zhang et al. (Zhang et al., 2022a) first decompose a GNN operation into four parts (i.e., scatter across edges, apply edge, gather from edges, and apply node), and then identify three challenges in GNN computations: redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. Specifically, redundant neural operators can be caused when the same feature transformation is employed to the same node features after they are scattered across edges. To address this issue, the scatter operation can be postponed after the feature transformation is finished. For the second issue, if the thread mappings are inconsistent (e.g., scatter threads are assigned with edge-centric tasks, while gather threads are assigned with node-centric tasks), these threads have to be processed sequentially and cannot be merged. This work leverages unified thread mapping for fusion. The third issue identifies that intermediate results may occupy up to 92% of the total memory requirement, which can be resolved by recomputation during backward propagation.
MaxK-GNN (Peng et al., 2024) accelerates GNN training on GPUs by addressing sparsity, irregularity, and memory inefficiencies in sparse matrix-matrix multiplication (SpMM) operations. It introduces a MaxK nonlinearity that induces controlled sparsity by selecting the top-k elements per node embedding, reducing effective feature dimensions with minimal accuracy loss while preserving universal approximation capabilities. To handle the resulting sparse features, it employs a Compressed Balanced Sparse Row (CBSR) format for contiguous storage and coalesced memory access. Custom kernels are designed, including a coalescing-enhanced forward SpGEMM (sparse general matrix-matrix multiplication) with shared memory-based accumulation and warp-level partitioning for workload balance, as well as an optimized backward SSpMM (sampled SpMM) with dense row prefetching and atomic global memory accumulation. These optimizations reduce global memory traffic significantly (up to 90%), improve cache hit rates, and enhance bandwidth utilization, leading to substantial speedups over baselines like cuSPARSE while maintaining comparable model accuracy.
D.2. User-defined Function Optimization
FeatGraph (Hu et al., 2020b) provides a flexible programming interface on top of the compiler stack for DL systems: TVM (Chen et al., 2018b) for supporting friendly user-defined functions. The users can manually select the tiling factor (the number of features being processed in each thread) for maximizing the cache utilization. However, because TVM lacks the support for sparse operations, FeatGraph only supports SpMM/SDDMM-based operations.
@@ D.2. User-defined Function Optimization
Graphiler (Xie et al., 2022) develops an automated tool for scatter/gather-related tasks. Specifically, compared with (Zhang et al., 2022a) mentioned in the previous section that manually implements operation reordering for some specific GNN operations, Graphiler enjoys better flexibility by automatically detecting redundant communication and then adjusting their order accordingly, and further fuses all operations after gathering from edges to reduce intermediate memory accesses.
gSampler (Gong et al., 2023) introduces a general and efficient GPU-based framework for graph sampling in graph learning tasks. It proposes the Extract-Compute-Select-Finalize (ECSF) programming model to unify a wide range of graph sampling algorithms, allowing users to express complex sampling logic through matrix-centric APIs. gSampler optimizes execution via a data flow graph intermediate representation, incorporating computation optimizations to eliminate redundant data movements, data layout optimizations to select efficient storage formats based on algorithms and datasets, and super-batching to enhance GPU utilization without affecting accuracy. Compared to systems like DGL, gSampler achieves average speedups of 6.54× in sampling, reducing overall GNN training time by over 40%.
D.3. On-device scalable training systems